Check whether the function f given by:

f(x)= (x-4)^2(x+1)^4 has local maxima and loca

minima


1
Expert's answer
2021-05-13T06:18:50-0400

Solution

f’(x) = 2(x-4) (x+1)4 +4(x-4)2(x+1)3 = 2(x-4) (x+1)3 [x+1+2(x-4)] = 2(x-4) (x+1)3 (3x-7)

So solutions of the equation f’(x) = 0 are x1 = -1, x2 = 7/3, x3 = 4. These points are points of extremum.

f’’(x) = 2 (x+1)3 (3x-7)+6(x-4) (x+1)2 (3x-7)+6(x-4) (x+1)3  = 2 (x+1)2 [(x+1)(3x-7)+3(x-4)(3x-7)+ 3(x-4) (x+1)] =  2 (x+1)2 [3x2-4x-7+3(x-4)(4x-6)] = 2 (x+1)2 [3x2-4x-7+12x2 -66x+72] = 2 (x+1)2 (15x2-70x+65)

f’’(x1) = 0, f’’(x2) = -370.37 <0, f’’(x3) = 1250>0.

Therefore x3=4 is a point of local minima, x2=7/3 is a point of local maxima

For point x1=-1 f(x1)=0 and for any ε>0 (x1± ε +1)4 >0 and f(x1± ε)>0. So x1=-1 is a point of local minima too.    

Answer

x1=-1 and x3=4 are points of local minima, x2=7/3 is a point of local maxima